OpenStudy (anonymous):

find all the angles between 0 and 360 degrees for tan2x = cot40. So we have the obvious method, with turning cot40 into a number, but is there any other method?? IB maths HL talking here :P

hartnn (hartnn):

yup, tan a = cot (90-a) so tan 2x = ?

OpenStudy (anonymous):

so 90 - 2x = 40 ??

hartnn (hartnn):

yup, x= ?

OpenStudy (anonymous):

x = 25. AHHH. MINDBLOWN :P

hartnn (hartnn):

one more solution is there

hartnn (hartnn):

tan(x+180) = tan x put x= 25 here

OpenStudy (anonymous):

of course, 205.

OpenStudy (anonymous):

hahahahahahah, thank youuu :) you have just saved me from the wrath of my math teacher :P

hartnn (hartnn):

welcome ^_^